OpenStudy (anonymous):

(-3/10,-2)is a solution to 10x-9y=15

OpenStudy (anonymous):

y yes it is

OpenStudy (anonymous):

Plug it in! If the values on the left equal the values on the right then it is a solution. Remember the x value is the first number in the parenthesis and the y value is the second number in the parenthesis.

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

show the work for this problem thank you

OpenStudy (anonymous):

(x,y) (-3/10,-2) 10x-9y=15 10(-3/10)-9(-2)=15 -3+18=15 15=15